OpenStudy (anonymous):

Simplify. 2 – 4 • 3 + 6 • 3

OpenStudy (anonymous):

answer is 8

OpenStudy (anonymous):

2-12+18

OpenStudy (igreen):

@panchajanya Don't give answers.

OpenStudy (anonymous):

i thought it was -28 because 12+18=30 then u subtract 2 from 30 since its on that side and u get -28

OpenStudy (igreen):

Well, you have to multiply 4 * 3 first.

OpenStudy (ahsome):

\(2−4×3+6×3\) Do the multiplying first, left to right \(2−4×3+6×3\) \(=2−12+6×3\) \(=2−12+18\) Now do the rest, left to right

OpenStudy (igreen):

That gives you: \(2 - 12 + 6 \cdot 3\) Now you multiply 6 * 3: \(2 - 12 + 18\) Now you do the rest, just subtract first, then add.

OpenStudy (anonymous):

so it is -28 right